SlideServe
Menu
Préférences
© SlideServe · v2026

Understanding Combinations: Calculating nCr for Different Scenarios

fiona avatar
40 diapositives · 683 Views · Aug 06, 2014

This guide explains combinations of n objects taken r at a time, where the order does not matter. The formula for combinations is nCr = n!/(r!(n-r)!), where n is the total number of objects and r is the number of objects chosen. For example, if 16 construction companies are bidding for a job, you can calculate the different combinations of 4 companies using this formula, resulting in 1820 unique combinations. This guide also includes a calculator method for quick determination of combinations.

Understanding Combinations: Calculating nCr for Different Scenarios - slide 1
01 / 40Understanding Combinations: Calculating nCr for Different Scenarios
1 / 17

Télécharger la présentation

Understanding Combinations: Calculating nCr for Different Scenarios
Une image / un lien ci-dessous est fourni (tel quel) pour télécharger la présentation

Politique de téléchargement : le contenu du site vous est fourni TEL QUEL pour votre information et votre usage personnel ; il ne peut être vendu, concédé sous licence ni partagé sur d’autres sites sans l’accord de son auteur.Le contenu vous est fourni TEL QUEL, pour votre information et votre usage personnel uniquement. Téléchargez la présentation en cliquant sur ce lien. Lors du téléchargement, si vous ne parvenez pas à télécharger une présentation, l’éditeur a peut-être supprimé le fichier de son serveur.Pendant le téléchargement, si vous ne parvenez pas à obtenir une présentation, le fichier a peut-être été supprimé par l’éditeur.

F I N

Plus de contenu associé

Transcription de la présentation
  1. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? .

  2. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? no particular order .

  3. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? no particular order . ₁₆C₄ =

  4. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? no particular order . ₁₆C₄ = 16!/(16-4)!4!

  5. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? no particular order . ₁₆C₄ = 16!/(16-4)!4! = 16!/(12!4!)

  6. 3.4C-Combinations • Combinationsof n objects taken r at a time • Order does NOT matter • nCr = n!/(n-r)!n! n=# objects r=# taken • Calculator: n, MATH, → PRB, 3:nCr, r, ENTER • Ex: 16 construction companies are bidding to do the job. How many different combinations of 4 companies can be chosen? no particular order . ₁₆C₄ = 16!/(16-4)!4! = 16!/(12!4!)=1820

  7. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  8. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  9. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  10. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  11. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  12. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) = 560 • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? .

  13. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) = 560 • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? order IS important (distinguishable permutation) .

  14. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) = 560 • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? order IS important (distinguishable permutation) . 11!/(1!4!4!2!)

  15. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) = 560 • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? order IS important (distinguishable permutation) . 11!/(1!4!4!2!) = 34650 possible & 1 favorable 1/34650

  16. More Examples: • There are 16 employees. In how many ways can an advisory committee of 3 be chosen? order NOT important ₁₆C₃ = 16!/(16-3)!3! = 16!/(13!3!) = 560 • A word has 1 M, 4 I’s, 4 S’s & 2 P’s. What is the probability they can be arranged to spell MISSISSIPPI? order IS important (distinguishable permutation) . 11!/(1!4!4!2!) = 34650 possible & 1 favorable 1/34650 = .000029

  17. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  18. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  19. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  20. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  21. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  22. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  23. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. .

  24. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important .

  25. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important P(desirable)/P(Total) .

  26. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important P(desirable)/P(Total)=P(5 of 13 D)/P(5 of 52) .

  27. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important P(desirable)/P(Total)=P(5 of 13 D)/P(5 of 52) . ₁₃C₅/₅₂C₅

  28. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important P(desirable)/P(Total)=P(5 of 13 D)/P(5 of 52) . ₁₃C₅/₅₂C₅ = 1287/2598960

  29. More Examples • A word has 1 L, 2 E’s, 2 T’s, & 1R. What is the probability they will be arranged to spell LETER? Order IS important Distinguishable permutation 6!/(1!2!2!1!) = 180 possible, 1 desirable = 1/180 = .0056 • Find the probability of being dealt 5 diamonds in a row from a standard deck of 52. Order NOT important P(desirable)/P(Total)=P(5 of 13 D)/P(5 of 52) . ₁₃C₅/₅₂C₅ = 1287/2598960 = .0005

  30. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men?

  31. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important

  32. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W)

  33. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ =

  34. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1

  35. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12)

  36. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12) = ₁₂C₃ =

  37. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12) = ₁₂C₃ = 220

  38. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12) = ₁₂C₃ = 220 P(all men) = P(3 M)·P( 0 W) / P(3 from 12)

  39. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12) = ₁₂C₃ = 220 P(all men) = P(3 M)·P( 0 W) / P(3 from 12) 10/220 =

  40. More Examples: • A group contains 5 men and 7 women. If 3 are selected at random, what is the probability that all 3 are men? Order is NOT important P(3 M)·P(0W) = ₅C₃· ₇C₀ = 10·1 =10 P(choose 3 from 12) = ₁₂C₃ = 220 P(all men) = P(3 M)·P( 0 W) / P(3 from 12) 10/220 = .045